# Believing without Seeing: Quality Scores for Contextualizing Vision-Language Model Explanations

### v1

@@ Appendix E Example of Model Explanations with Different Qualities
From Table˜8, we show our evaluation framework on four representative VQA instances. For each example, we provide the original question (and answer choices where applicable), image, ground truth answer, the model’s selected answer and rationale, and the corresponding scores for Simulatability, Informativeness, Commonsense Plausibility, Visual Fidelity, and Contrastiveness. Note that for VizWiz, the ground-truth answer is defined as the majority vote among the 10 user annotations from the original dataset, and since VizWiz provides no answer choices, we do not compute a Contrastiveness score.
Table 8: Four VQA examples from both A-OKVQA and VizWiz datasets with model predictions, explanations, and qualities. The predicted answer and explanation are from LLaVA-v1.5-7B model.
@@
Dataset Question Answer Choices Image Ground Truth Predicted Answer LLaVA-v1.5-7B Model Explanation Sim. Info. Plau. VF Contr.
@@
VizWiz What color are those pants? [Uncaptioned image] pink pink The man is wearing pink pajamas, which is an unusual choice for a casual setting. 0.025 1.000 0.397 1.000
@@
VizWiz What is in this bottle? [Uncaptioned image] bacon bits spices The person holding the bottle of spices is likely using it to season their food, adding flavor and aroma to various dishes. Spices are commonly used in cooking as they can enhance the taste of ingredients and create unique flavor combinations. The hand holding the bottle suggests that the person might be preparing a meal or experimenting with different spice blends for their culinary creations. 0.999 1.000 0.566 0.500
Appendix F Reliability and Validity of the Trust Cues
F.1 Validation on Agreement with Human Judgements
@@
To validate the reliability of our automated Visual Fidelity and Contrastiveness qualities, we conducted an internal sanity check on a held‐out sample. We randomly selected 100 examples from our evaluation sets and had a single expert annotator manually assess each instance. The instruction provided to the expert annotator is shown at Table˜11. We binarized our scores at a threshold of $0.5$ (scores $\geq 0.5$ map to $1$, else 0) and compared these labels to the human judgments using Cohen’s $\kappa$:
@@
F.2 Accuracy of LLM-generated verification questions and answers
@@
F.3 Verifier robustness for Visual Fidelity
@@
As shown in Table˜9 and Table˜10, across A-OKVQA and VizWiz, all three verifiers yield positive discriminability on average (GPT-4o: 0.109; Gemma-3n-E4B: 0.061; Qwen2.5-VL-7B: 0.154), and ECE remains comparably low (0.168/0.197/0.196 respectively). Thus, VF continues to separate correct from incorrect predictions and stays well-calibrated even with using smaller open models as verifiers.
Table 11: Instructions Provided to the Expert Annotator
@@
IMPORTANT: See examples from this form before proceeding!
@@
IMPORTANT: See examples from this form before proceeding!
Appendix G User Studies
G.1 Annotator filtering
Table 12: Prolific annotator filtering
@@ G.1 Annotator filtering
To ensure high-quality annotations, we recruited annotators on Prolific using the criteria listed in Table˜12. All participants were required to be native English speakers residing in the United States, with approval rates between 98–100% and at least 1000 prior submissions. We further restricted enrollment to individuals holding at least an undergraduate degree. To prevent contamination across experimental conditions, each participant was confined to a single “setting” (i.e., one quality-type configuration).
G.2 Creating the subset splits
@@
For our qualities evaluations from Section˜3, we selected two 500-instance sets. From A-OKVQA we used the first 500 examples from the official validation split. From VizWiz we selected the first 500 validation examples whose majority human annotation was not “unanswerable” and which contained no NSFW content, as filtered by GPT-4o.
To create the subset used for human studies, we create a subset of 100 questions each from the A-OKVQA and VizWiz 500-instance datasets above. These subsets are chosen by sampling 50 times from the validation set; each sampled subset contains randomly sampled 50 correct and 50 incorrect instances from the validation set. On the A-OKVQA dataset, we picked the subset for which the average Expected Calibration Error across Visual Fidelity and Contrastiveness is the lowest in 50 samples; on the VizWiz dataset, we picked the subset for which the Expected Calibration Error for Visual Fidelity is the lowest in our 50 samples, as Contrastiveness is not available on the VizWiz dataset. From Table˜13 this maintains a similar ECE distribution of these quality measures to the full set from Section˜3.
@@
Table 13: Accuracy and Expected Calibration Error (ECE) for different qualities. “Subset selected” rows correspond to the 100-question subsets (50 correct / 50 incorrect) selected and the other two rows represent LLaVA-v1.5-7B on the A-OKVQA and VizWiz 500-instance datasets.
[…]
AOKVQA 0.696 0.207 0.176 0.164 0.147 0.109 0.288 0.332 0.162
AOKVQA (subset selected) 0.500 0.270 0.237 0.133 0.130 0.227 0.372 0.490 0.075
VizWiz 0.557 0.271 0.305 0.429 0.110
VizWiz (subset selected) 0.500 0.236 0.379 0.390 0.071
G.3 Attention Incentives
@@ G.3.2 Bonus Payments
G.4 User Studies Settings & Examples
Table 14: Summary of the 14 one‐step human‐study settings.
@@ G.4 User Studies Settings & Examples
Show Explanation Only No quality scores shown
@@
[…]
@@
VF num Show Visual Fidelity numeric score “AI Confidence that the explanation accurately describes the image details”
VF desc Show at most two descriptive sentences converted from visual questions which are verified by verifier VLM $m_{\texttt{Verif}}$, and at most two from questions which are not verified by $m_{\texttt{Verif}}$
Contr. num Show Contrastiveness score “AI Confidence that the explanation rules out the other choices”
Contr. desc Show the other answer options $a_{j}\neq a_{0}$ s.t. $\textsc{Pr}_{\texttt{NLI}}(P\text{ entails }h_{j})\geq 0.5$
Both Numeric Display both VF num and Contr. num messages
Both Descriptive Display both VF desc and Contr. desc messages
VF shown as Conf VF score displayed as a simple confidence score “AI Confidence that the explanation is correct”
Contr shown as Conf Contr. score displayed as a simple confidence score “AI Confidence that the explanation is correct”
@@
Table 15: User-study results on A-OKVQA and VizWiz in the one stage setting. Each row received 300 annotations (100 questions $\times$ 3 annotations per question).
@@
[…]
AVG(VF, Contr.) 7.3% ± 1.5% 74.8% ± 2.6% 66.2% ± 2.8% 52.7% ± 2.5% 10.0% ± 1.3%
VF num 7.3% ± 1.5% 76.3% ± 2.6% 62.9% ± 2.9% 58.7% ± 2.6% 10.0% ± 1.3%
VF desc 9.7% ± 1.7% 69.4% ± 2.8% 68.6% ± 2.8% 45.3% ± 2.4% 11.3% ± 1.3%
Contr. num 9.0% ± 1.7% 77.7% ± 2.5% 64.1% ± 2.9% 56.7% ± 2.6% 8.7% ± 1.2%
Contr. desc 6.0% ± 1.4% 71.6% ± 2.7% 61.3% ± 2.9% 56.7% ± 2.6% 16.0% ± 1.6%
Both Numeric 10.7% ± 1.8% 72.8% ± 2.7% 66.4% ± 2.9% 48.7% ± 2.5% 11.3% ± 1.3%
Both Descriptive 8.7% ± 1.6% 68.2% ± 2.8% 67.2% ± 2.8% 46.0% ± 2.4% 14.0% ± 1.5%
VF shown as Conf 7.0% ± 1.5% 77.8% ± 2.5% 61.3% ± 2.9% 60.0% ± 2.7% 12.0% ± 1.4%
Contr. shown as Conf 8.3% ± 1.6% 74.5% ± 2.6% 69.1% ± 2.8% 50.0% ± 2.5% 6.7% ± 1.0%
Prod shown as VF 7.7% ± 1.5% 65.7% ± 2.9% 67.5% ± 2.8% 44.4% ± 2.4% 16.0% ± 1.6%
Prod shown as Contr. 8.0% ± 1.6% 70.3% ± 2.8% 68.5% ± 2.8% 44.7% ± 2.4% 13.3% ± 1.4%
VizWiz Control 7.3% ± 1.5% 70.5% ± 2.7% 59.4% ± 3.0% 55.3% ± 2.6% 20.0% ± 1.7%
VF num 9.3% ± 1.7% 65.1% ± 2.9% 64.7% ± 2.9% 44.7% ± 2.4% 19.3% ± 1.7%
VF desc 10.3% ± 1.8% 67.3% ± 2.9% 69.1% ± 2.8% 42.0% ± 2.4% 13.3% ± 1.4%
In the main user study, we present 14 settings in total. Table˜14 summarizes all 14 different settings and Figure˜9 shows examples of all 8 different categories of messages presented to users. Table˜15 contains the main user-study results of user behavior patterns.
G.5 Supplementary Human Studies
In addition to the study described in §4.2 and Section˜G.4, we ran a supplementary user study to evaluate how explanations and qualities affect users’ trust in model predictions. Each of the 10 questions was presented in three successive stages; after each stage, participants indicated whether they believed the model’s answer was correct or were unsure:
@@ G.5 Supplementary Human Studies
Participants were paid a \$2 base fee and could earn up to \$1 in performance‐based bonuses, which were awarded only during Stage 3 (Explanation + Quality; see Section˜G.3.2).
Table 16: User‐study results on A‑OKVQA and VizWiz in the three stages setting.
@@
A‑OKVQA LLaVA-v1.5-7B VF (numeric) 300 $65.3\%\pm 2.8\%$ $67.3\%\pm 4.6\%$ $19.0\%\pm 2.3\%$ $63.4\%\pm 3.1\%$ $5.0\%\pm 1.3\%$ $63.5\%\pm 2.9\%$
Contr (numeric) 300 $65.7\%\pm 2.7\%$ $58.3\%\pm 4.9\%$ $17.7\%\pm 2.2\%$ $60.7\%\pm 3.1\%$ $6.3\%\pm 1.4\%$ $61.2\%\pm 2.9\%$
Both VF and Contr (numeric) 300 $77.3\%\pm 2.4\%$ $58.8\%\pm 6.0\%$ $38.0\%\pm 2.8\%$ $54.8\%\pm 3.7\%$ $9.7\%\pm 1.7\%$ $64.2\%\pm 2.9\%$
Avg(VF, Contr) 300 $86.0\%\pm 2.0\%$ $54.8\%\pm 7.8\%$ $35.3\%\pm 2.8\%$ $63.9\%\pm 3.5\%$ $6.0\%\pm 1.4\%$ $63.5\%\pm 2.9\%$
VF (descriptive) 300 $67.3\%\pm 2.7\%$ $59.2\%\pm 5.0\%$ $19.7\%\pm 2.3\%$ $59.3\%\pm 3.2\%$ $3.7\%\pm 1.1\%$ $62.3\%\pm 2.9\%$
Contr (descriptive) 300 $63.0\%\pm 2.8\%$ $64.9\%\pm 4.6\%$ $17.0\%\pm 2.2\%$ $63.1\%\pm 3.1\%$ $8.7\%\pm 1.6\%$ $59.5\%\pm 3.0\%$
Both VF and Contr (descriptive) 290 $70.7\%\pm 2.7\%$ $47.1\%\pm 5.4\%$ $23.8\%\pm 2.5\%$ $57.9\%\pm 3.3\%$ $8.3\%\pm 1.6\%$ $64.7\%\pm 2.9\%$
VizWiz Qwen2.5-VL-7B VF (numeric) 300 $57.7\%\pm 2.9\%$ $59.8\%\pm 4.4\%$ $5.7\%\pm 1.3\%$ $65.7\%\pm 2.8\%$ $1.7\%\pm 0.7\%$ $65.1\%\pm 2.8\%$
VF (descriptive) 300 $58.7\%\pm 2.8\%$ $66.1\%\pm 4.3\%$ $17.7\%\pm 2.2\%$ $65.2\%\pm 3.0\%$ $4.7\%\pm 1.2\%$ $67.5\%\pm 2.8\%$
Table˜16 shows that as users progress from seeing only the model’s answer to viewing explanations and then explanations with quality scores, their unsure rate steadily decreases while user accuracy correspondingly increases, with the largest gains observed when descriptive qualities are provided alongside explanations. These findings support our findings that richer, more interpretable quality signals can meaningfully improve users’ trust calibration.
Appendix H Computational Resources Spent & Total Cost
The complete breakdown of all monetary expenditures across the study is given in Table˜17. In brief, the computational expenses include six hours of A100 GPU time for running the vision‐language models, plus OpenAI API calls for generating predicted answers, explanations, and computing qualities. Human evaluation costs cover two Prolific studies—a one‐stage study as shown in Section˜G.4 and a three‐stage follow‐up from Section˜G.5 (including up to \$1.00 in per‐participation bonuses; see Section˜G.3.2).

### v2

@@ Appendix H Uncertainty Quantification Baselines
From Table˜13, we show our evaluation framework on four representative VQA instances. For each example, we provide the original question (and answer choices where applicable), image, ground truth answer, the model’s selected answer and rationale, and the corresponding scores for Simulatability, Informativeness, Commonsense Plausibility, Visual Fidelity, and Contrastiveness. Note that for VizWiz, the ground-truth answer is defined as the majority vote among the 10 user annotations from the original dataset, and since VizWiz provides no answer choices, we do not compute a Contrastiveness score.
@@
Dataset Question Answer Choices Image Ground Truth Predicted Answer LLaVA-v1.5-7B Explanation Sim. Info. Plau. VF Contr.
@@
VizWiz What color are those pants? blue, red, green [Uncaptioned image] pink, pink white checkered pink The man is wearing pink pajamas, which is an unusual choice for a casual setting. 0.025 1.000 0.397 1.000 0.073
@@
VizWiz What is in this bottle? pen, laser pointer, small torch [Uncaptioned image] bacon bits, hormel bacon bits, bacon spices The person holding the bottle of spices is likely using it to season their food, adding flavor and aroma to various dishes. Spices are commonly used in cooking as they can enhance the taste of ingredients and create unique flavor combinations. The hand holding the bottle suggests that the person might be preparing a meal or experimenting with different spice blends for their culinary creations. 0.999 1.000 0.566 0.500 0.555
Table 13: Four VQA examples from both A-OKVQA and VizWiz datasets with model predictions, explanations, and qualities. The predicted answer and explanation are from LLaVA-v1.5-7B model. Answer choices in VizWiz examples (shown in gray) are generated negatives from the ground truth, which are not shown to the VLM model when generating the predictions and explanations and are only provided to compute the contrastiveness (shown in gray).
Appendix I VizWiz Dataset Limitations and Negative Discriminability
We observed that for VizWiz, particularly with GPT-4o, the discriminability scores for our proposed metrics (and baselines) are often low or negative (see Table˜1). A negative discriminability score implies that the model assigns higher quality scores to “incorrect” predictions than to “correct” ones.
Upon qualitative inspection, we found this phenomenon is largely driven by the “Ground Truth Problem” inherent in the VizWiz dataset. The ground truth answers are often narrow or specific, causing valid model predictions to be labeled as incorrect. Table˜14 shows three representative examples:
Question Image Ground Truth List Model Prediction Explanation VF Contr.
What does the screen say? [Uncaptioned image] windows 7 home premium, windows, windows home premium GPT-4o set up windows The screen displays the setup process for Windows 7 Home Premium. It prompts the user to choose a user name for their account and name their computer to distinguish it on the network. The text on the screen includes: - "Set Up Windows" - "Windows 7 Home Premium" - "Choose a user name for your account and name your computer to distinguish it on the network." - "Type a user name (for example, John):" - "Type a computer name:" - "PC" - "Copyright © 2009 Microsoft Corporation. All rights reserved." 1.000 0.585
What is this? [Uncaptioned image] tall buildings, buildings, city landscape, downtown city, city buildings, skyscrapers, city GPT-4o downtown Toronto The image shows a cityscape with several tall buildings, one of which has the TD (Toronto-Dominion) logo prominently displayed. This suggests that the location is likely in downtown Toronto, where TD Bank has a significant presence. The mix of modern and older architecture is also characteristic of downtown Toronto. Therefore, the answer is indeed downtown Toronto. 1.000 0.659
What’s this? [Uncaptioned image] room, half computer screen, monitor, laptop, computer screen, computer monitor, tv Qwen2.5-VL-7B keyboard The image shows a piece of electronic equipment that appears to be an electronic keyboard or synthesizer. It has keys similar to a piano and is positioned next to what looks like a computer monitor, suggesting it might be used for music production or performance. The lighting is dim, and the image quality is low, making it difficult to discern finer details, but the shape and layout of the keys are consistent with a keyboard. 1.000 0.999
Table 14: Examples of “Incorrect” VizWiz predictions that received high Explanation Quality scores. In these cases, the model provides a semantically valid or more specific answer than the ground truth, leading to a “False Negative” error judgment despite high Visual Fidelity and Contrastiveness.
  • Setup Windows Example: The ground truth includes “windows 7 home premium”, but the GPT-4o model predicts “set up windows”, which accurately describes the visible title. This is marked incorrect, yet receives high quality scores.
  • Toronto Example: The ground truth is generic (“tall buildings”), but the GPT-4o model correctly identifies “downtown Toronto”. This specific identification is penalized, again resulting in high quality scores for a “wrong” answer.
  • Keyboard Example: The ground truth lists various terms for a screen (e.g., “monitor”, “laptop”), but Qwen2.5-VL-7B predicts “keyboard”, identifying the other prominent object in the blurry image. The explanation correctly describes the keyboard’s features, leading to high fidelity despite the answer mismatch.
To quantify this, an expert annotator manually inspected 20 randomly sampled instances where the model prediction was marked “incorrect” but achieved a Visual Fidelity score >0.9>0.9. We found that for GPT-4o, 60% (12/20) of these cases were actually valid answers rejected by the ground truth (e.g., specific vs generic). This rate was lower for Qwen2.5-VL-7B (25%, 5/20) and LLaVA-v1.5-7B (15%, 3/20), confirming that GPT-4o’s “hallucinations” on VizWiz are frequently artifacts of valid but mismatched reasoning.
Because GPT-4o often provides these “over-answering” or semantically valid but lexically distinct responses, it accumulates a large number of high-quality “incorrect” samples, which inverts the discriminability metric. This effect is less pronounced in Qwen2.5-VL-7B and least in LLaVA-v1.5-7B, which tend to generate shorter, simpler answers closer to the crowd-sourced ground truths. Thus, the negative discriminability reflects a limitation of the evaluation ground truth rather than a failure of the explanation quality metrics.
Appendix J Reliability and Validity of the Trust Cues
J.1 Validation on Agreement with Human Judgements
@@ Appendix I VizWiz Dataset Limitations and Negative Discriminability
To validate the reliability of our automated Visual Fidelity and Contrastiveness qualities, we conducted an internal sanity check on a held‐out sample. We randomly selected 100 examples from our evaluation sets and had a single expert annotator manually assess each instance. The instruction provided to the expert annotator is shown at Table˜17. We binarized our scores at a threshold of $0.5$ (scores $\geq 0.5$ map to $1$, else 0) and compared these labels to the human judgments using Cohen’s $\kappa$:
@@
J.2 Accuracy of LLM-generated verification questions and answers
@@
J.3 Verifier robustness for Visual Fidelity
@@
As shown in Table˜15 and Table˜16, across A-OKVQA and VizWiz, all three verifiers yield positive discriminability on average (GPT-4o: 0.109; Gemma-3n-E4B: 0.061; Qwen2.5-VL-7B: 0.154), and ECE remains comparably low (0.168/0.197/0.196 respectively). Thus, VF continues to separate correct from incorrect predictions and stays well-calibrated even with using smaller open models as verifiers.
@@
IMPORTANT: See examples from this form (https://forms.gle/nvpcQ7fpLoFqsLiJ9) before proceeding!
@@
IMPORTANT: See examples from this form (https://forms.gle/HYp6N44TpsjAoxHa6) before proceeding!
Table 17: Instructions Provided to the Expert Annotator
Appendix K User Studies
K.1 Annotator filtering
@@
Table 18: Prolific annotator filtering
To ensure high-quality annotations, we recruited annotators on Prolific using the criteria listed in Table˜18. All participants identities were fully anonymized. All annotators provided informed consent through Prolific’s platform, which clearly described the task purpose, data use, and compensation prior to participation.
This study protocol was reviewed and approved by our institution’s Institutional Review Board (the name of the institution will be included in the camera-ready version).
@@
All participants were required to be native English speakers residing in the United States, with approval rates between 98–100% and at least 1000 prior submissions. We further restricted enrollment to individuals holding at least an undergraduate degree. To prevent contamination across experimental conditions, each participant was confined to a single “setting” (i.e., one quality-type configuration).
K.2 Dataset Splits and User Study Design
Metric Evaluation Sets (500 instances)
@@
For the quality evaluations detailed in Section˜3, we selected three 500-instance sets. From A-OKVQA, we used the first 500 examples from the official validation split. From VizWiz, we selected the first 500 validation examples where the majority human annotation was not “unanswerable” and which contained no NSFW content, as filtered by GPT-4o. From MMMU-Pro, we used the first 500 examples with single-image contexts from the standard (10 options) set.
Human Study Subsets (100 instances)
We conducted a between-subjects user study where participants predicted VLM correctness based on varying levels of explanation quality scores. To facilitate this, we derived a subset of 100 questions each from the A-OKVQA and VizWiz evaluation sets described above.
These subsets were balanced to contain 50 correct and 50 incorrect predictions by LLaVA-v1.5-7B. To ensure the subsets remained representative of the larger distribution, we generated 50 random candidate subsets and selected the one with the lowest Expected Calibration Error (ECE) relative to the full 500-instance set. Specifically, for A-OKVQA, we minimized the average ECE across Visual Fidelity and Contrastiveness; for VizWiz, we minimized the ECE for Visual Fidelity (as Contrastiveness scores were unavailable during the subset creation phase). As shown in Table˜20 and Table˜20, this selection maintains a distribution of quality scores consistent with the full set from Section˜3.
Study Protocol
In the study, each question was annotated by three unique participants, yielding 300 annotations per dataset per condition. Each participant completed a batch of 10 questions, balanced to include five instances where the model was correct and five where it was incorrect; participants were not informed of this distribution. Users were randomly assigned to a single study condition (control or a treatment variant) and did not participate in multiple conditions. In total, we collect 300 human annotations per dataset and condition (3 users × 100 questions), capturing user reliance behavior under varying explanation and quality signal conditions.
Set Accuracy VF Contr. Prod(VF, Contr.) Min(VF, Contr.) Avg(VF, Contr.) Support Informative Plausibility
Full (500) 0.696 0.207 0.176 0.164 0.147 0.109 0.288 0.332 0.162
Subset (100) 0.500 0.270 0.237 0.133 0.130 0.227 0.372 0.490 0.075
@@
Table 19: Accuracy and Expected Calibration Error (ECE) for different qualities on the A-OKVQA dataset. “Subset” rows correspond to the 100-question user study subsets (50 correct / 50 incorrect), while “Full” rows represent the 500-instance evaluation set.
Set Accuracy VF Support Informative Plausibility
Full (500) 0.557 0.271 0.305 0.429 0.110
Subset (100) 0.500 0.236 0.379 0.390 0.071
Table 20: Accuracy and Expected Calibration Error (ECE) for different qualities on the VizWiz dataset.
K.3 Attention Incentives
@@
K.4 User Studies Settings & Examples
@@
Control Show explanation only; no quality scores are shown.
@@
Prod(VF, Contr) Product of VF and Contrast scores, shown as a simple confidence score “AI Confidence that the explanation is correct”.
@@
VF num. Show Visual Fidelity numeric score “AI Confidence that the explanation accurately describes the image details”
VF desc. Show at most two descriptive sentences converted from visual questions which are verified by verifier VLM $m_{\texttt{Verif}}$, and at most two from questions which are not verified by $m_{\texttt{Verif}}$
Contr. num. Show Contrastiveness score “AI Confidence that the explanation rules out the other choices”
Contr. desc. Show the other answer options $a_{j}\neq a_{0}$ s.t. $\textsc{Pr}_{\texttt{NLI}}(P\text{ entails }h_{j})\geq 0.5$
Both num. Display both VF num. and Contr. num. messages
Both desc. Display both VF desc. and Contr. desc. messages
VF shown as Conf. VF score displayed as a simple confidence score “AI Confidence that the explanation is correct”
Contr shown as Conf. Contr. score displayed as a simple confidence score “AI Confidence that the explanation is correct”
@@
Table 21: Summary of the 14 one‐step human‐study settings.
@@
Prod(VF, Contr.) 09.3% ± 1.7% 69.9% ± 2.8% 70.2% ± 2.8% 45.3% ± 2.4% 08.7% ± 1.2%
Avg(VF, Contr.) 07.3% ± 1.5% 74.8% ± 2.6% 66.2% ± 2.8% 52.7% ± 2.5% 10.0% ± 1.3%
VF num. 07.3% ± 1.5% 76.3% ± 2.6% 62.9% ± 2.9% 58.7% ± 2.6% 10.0% ± 1.3%
VF desc. 09.7% ± 1.7% 69.4% ± 2.8% 68.6% ± 2.8% 45.3% ± 2.4% 11.3% ± 1.3%
Contr. num. 09.0% ± 1.7% 77.7% ± 2.5% 64.1% ± 2.9% 56.7% ± 2.6% 08.7% ± 1.2%
Contr. desc. 06.0% ± 1.4% 71.6% ± 2.7% 61.3% ± 2.9% 56.7% ± 2.6% 16.0% ± 1.6%
Both num. 10.7% ± 1.8% 72.8% ± 2.7% 66.4% ± 2.9% 48.7% ± 2.5% 11.3% ± 1.3%
Both desc. 08.7% ± 1.6% 68.2% ± 2.8% 67.2% ± 2.8% 46.0% ± 2.4% 14.0% ± 1.5%
VF shown as Conf. 07.0% ± 1.5% 77.8% ± 2.5% 61.3% ± 2.9% 60.0% ± 2.7% 12.0% ± 1.4%
Contr. shown as Conf. 08.3% ± 1.6% 74.5% ± 2.6% 69.1% ± 2.8% 50.0% ± 2.5% 06.7% ± 1.0%
Prod shown as VF 07.7% ± 1.5% 65.7% ± 2.9% 67.5% ± 2.8% 44.4% ± 2.4% 16.0% ± 1.6%
Prod shown as Contr. 08.0% ± 1.6% 70.3% ± 2.8% 68.5% ± 2.8% 44.7% ± 2.4% 13.3% ± 1.4%
VizWiz Control 07.3% ± 1.5% 70.5% ± 2.7% 59.4% ± 3.0% 55.3% ± 2.6% 20.0% ± 1.7%
VF num. 09.3% ± 1.7% 65.1% ± 2.9% 64.7% ± 2.9% 44.7% ± 2.4% 19.3% ± 1.7%
VF desc. 10.3% ± 1.8% 67.3% ± 2.9% 69.1% ± 2.8% 42.0% ± 2.4% 13.3% ± 1.4%
Table 22: User-study results on A-OKVQA and VizWiz in the one-stage setting. Each row received 300 annotations (100 questions $\times$ 3 annotations per question).
In the main user study, we present 14 settings in total. Table˜21 summarizes all 14 different settings and Figure˜9 shows examples of all 8 different categories of messages presented to users. Table˜22 contains the main user-study results of user behavior patterns.
K.5 Supplementary Human Studies
In addition to the study described in Section˜4.2 and Section˜K.4, we ran a supplementary user study to evaluate how explanations and qualities affect users’ trust in model predictions. Each of the 10 questions was presented in three successive stages; after each stage, participants indicated whether they believed the model’s answer was correct or were unsure:
@@
Participants were paid a \$2 base fee and could earn up to \$1 in performance‐based bonuses, which were awarded only during Stage 3 (Explanation + Quality; see Section˜K.3.2).
@@
A‑OKVQA LLaVA VF num. 300 $65.3\%\pm 2.8\%$ $67.3\%\pm 4.6\%$ $19.0\%\pm 2.3\%$ $63.4\%\pm 3.1\%$ $5.0\%\pm 1.3\%$ $63.5\%\pm 2.9\%$
Contr num. 300 $65.7\%\pm 2.7\%$ $58.3\%\pm 4.9\%$ $17.7\%\pm 2.2\%$ $60.7\%\pm 3.1\%$ $6.3\%\pm 1.4\%$ $61.2\%\pm 2.9\%$
Both num. 300 $77.3\%\pm 2.4\%$ $58.8\%\pm 6.0\%$ $38.0\%\pm 2.8\%$ $54.8\%\pm 3.7\%$ $9.7\%\pm 1.7\%$ $64.2\%\pm 2.9\%$
Avg(VF, Contr.) 300 $86.0\%\pm 2.0\%$ $54.8\%\pm 7.8\%$ $35.3\%\pm 2.8\%$ $63.9\%\pm 3.5\%$ $6.0\%\pm 1.4\%$ $63.5\%\pm 2.9\%$
VF desc. 300 $67.3\%\pm 2.7\%$ $59.2\%\pm 5.0\%$ $19.7\%\pm 2.3\%$ $59.3\%\pm 3.2\%$ $3.7\%\pm 1.1\%$ $62.3\%\pm 2.9\%$
Contr desc. 300 $63.0\%\pm 2.8\%$ $64.9\%\pm 4.6\%$ $17.0\%\pm 2.2\%$ $63.1\%\pm 3.1\%$ $8.7\%\pm 1.6\%$ $59.5\%\pm 3.0\%$
Both desc. 290 $70.7\%\pm 2.7\%$ $47.1\%\pm 5.4\%$ $23.8\%\pm 2.5\%$ $57.9\%\pm 3.3\%$ $8.3\%\pm 1.6\%$ $64.7\%\pm 2.9\%$
VizWiz Qwen2.5 VF num. 300 $57.7\%\pm 2.9\%$ $59.8\%\pm 4.4\%$ $05.7\%\pm 1.3\%$ $65.7\%\pm 2.8\%$ $1.7\%\pm 0.7\%$ $65.1\%\pm 2.8\%$
VF desc. 300 $58.7\%\pm 2.8\%$ $66.1\%\pm 4.3\%$ $17.7\%\pm 2.2\%$ $65.2\%\pm 3.0\%$ $4.7\%\pm 1.2\%$ $67.5\%\pm 2.8\%$
Table 23: User-study results on A‑OKVQA and VizWiz in the three-stages setting.
Table˜23 shows that as users progress from seeing only the model’s answer to viewing explanations and then explanations with quality scores, their unsure rate steadily decreases while user accuracy correspondingly increases, with the largest gains observed when descriptive qualities are provided alongside explanations. These findings support our findings that richer, more interpretable quality signals can meaningfully improve users’ trust calibration.
Appendix L Computational Resources Spent & Total Cost
The complete breakdown of all monetary expenditures across the study is given in Table˜24. In brief, the computational expenses include six hours of A100 GPU time for running the vision‐language models, plus OpenAI API calls for generating predicted answers, explanations, and computing qualities. Human evaluation costs cover two Prolific studies—a one‐stage study as shown in Section˜K.4 and a three‐stage follow‐up from Section˜K.5 (including up to \$1.00 in per‐participation bonuses; see Section˜K.3.2).